If the angles of a triangle are in the ratio 2:3:7, the triangle is?
Answer by jim_thompson5910(35256)

2x+3x+7x = 180

12x = 180

x = 180/12

x = 15

2x = 2*15 = 30

3x = 3*15 = 45

7x = 7*15 = 105


So the three angles are 30, 45 and 105 degrees
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------